40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Oct 1 17:47:21 2011, Koji, Summary, General, Recovery from the power shutdown 
    Reply  Sat Oct 1 22:53:49 2011, kiwamu, Summary, General, Recovery from the power shutdown : lockable 
    Reply  Sun Oct 2 13:45:13 2011, rana, Summary, General, Recovery from the power shutdown: apache / svn 
    Reply  Mon Oct 3 08:38:21 2011, steve, Summary, VAC, Recovery from the power shutdown is completed poweroutage.jpg
       Reply  Thu Oct 13 21:40:59 2011, rana, Summary, VAC, Recovery from the power shutdown is completed 
Message ID: 5593     Entry time: Sat Oct 1 22:53:49 2011     In reply to: 5592
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Recovery from the power shutdown : lockable 

Found the Marconi for the 11 MHz source had been reset to its default.

 => reset the parameters. f = 11.065910 MHz (see #5530) amp = 13 dBm.

Interferometer became lockable. I checked the X/Y arm lock, and MICH lock, they all are fine.

ELOG V3.1.3-